### v1

# Extremality of families of sets and set-valued optimization

Abstract
The paper explores a new extremality model involving collections of arbitrary families of sets. We demonstrate its applicability to set-valued optimization problems with general preferences, weakening the assumptions of the known results and streamlining their proofs.
keywords:
extremal principle; separation; optimality conditions; set-valued optimization
{amscode}
49J52; 49J53; 49K40; 90C30; 90C46
1 Introduction
The paper explores another extension of the extremal principle [1, 2, 3].
Recall that a Banach space is Asplund if every continuous convex function on an open convex set is Fréchet differentiable on a dense subset [4], or equivalently, if the dual of each its separable subspace is separable. We refer the reader to [4, 3] for discussions about and characterizations of Asplund spaces. All reflexive, particularly, all finite dimensional Banach spaces are Asplund.
The core feature of the definitions of extremality and stationarity is the non-intersection of certain perturbations of either the original sets $\Omega_{1},\ldots,\Omega_{n}$ or their translations $\Omega_{1}-x_{1},\ldots,\Omega_{n}-x_{n}$. The perturbations employed in the definitions are in the form of linear translations of the sets: given a set $\Omega$, its linear translation is the set $\widetilde{\Omega}:=\Omega-a$ for some vector $a$.
The conventional extremal principle has proved to be versatile enough to cover a wide range of problems in optimization and variational analysis as demonstrated, e.g., in the books [5, 3, 6]. At the same time, there exist problems, mainly in multiobjective and set-valued optimization described via closed preference relations, that cannot be covered within the framework of linear translations. The first example of this kind was identified in Zhu [7]. Fortunately, such problems can be handled with the help of a more flexible extended version of the extremal principle using more general nonlinear perturbations (deformations) of the sets defined by set-valued mappings. Such an extension was developed in Mordukhovich et al. [8] (see also [5, 6]) and applied by other authors to various multiobjective problems in [9, 10, 11].
The model in [8] exploits non-intersection of perturbations (deformations) of the given sets $\Omega_{1},\ldots,\Omega_{n}$. The perturbations are chosen from the respective families of sets $\Xi_{i}:=\{S_{i}(s)\mid s\in M_{i}\}$ $(i=1,\ldots,n)$ determined by the given set-valued mappings $S_{1},\ldots,S_{n}$. In the particular case of linear translations, i.e. when, for all $i=1,\ldots,n$, $(M_{i},d_{i})=(X,d)$ and $S_{i}(a)=\Omega_{i}-a_{i}$ $(a_{i}\in X)$, we have $S_{i}(0)=\Omega_{i}$, and the property reduces to the one in the conventional extremal principle. It was shown by examples in [8, 6] that in general the framework of set-valued perturbations is richer than that of linear translations. With minor modifications in the proof, the conventional extremal principle can be extended to the set-valued setting producing a more advanced version of fuzzy separation; cf. [8, 6].
The extremality model based on set-valued perturbations has proved to be useful when dealing with multiobjective optimization problems with general preference relations. This model has been revisited and refined in [12], making it more flexible and, at the same time, simpler. The refined model studies extremality of nonempty families of arbitrary sets $\Xi_{i}$ $(i=1,\ldots,n)$ and is applicable to a wider range of variational problems. The next definition and theorem are taken from [12, Definition 3.2 and Theorem 3.4], respectively.
Definition 1.1 (Extremality and stationarity: families of sets).
Let $\Xi_{1},\ldots,\Xi_{n}$ be families of subsets of a normed space $X$, and $\bar{x}\in X$.
  1. i.
    The collection $\{\Xi_{1},\ldots,\Xi_{n}\}$ is extremal at $\bar{x}$ if there is a $\rho\in(0,+\infty]$ such that, for any $\varepsilon>0$, there exist $A_{i}\in\Xi_{i}$ $(i=1,\ldots,n)$ such that $\max_{1\leq i\leq n}d(\bar{x},A_{i})<\varepsilon$ and $\bigcap_{i=1}^{n}A_{i}\cap B_{\rho}(\bar{x})=\emptyset$.
  2. ii.
    Let $\Omega_{1},\ldots,\Omega_{n}\subset X$. The collection $\{\Xi_{1},\ldots,\Xi_{n}\}$ is approximately stationary at $\bar{x}$ with respect to $\{\Omega_{1},\ldots,\Omega_{n}\}$ if, for any $\varepsilon>0$, there exist a $\rho\in(0,\varepsilon)$, $x_{i}\in\Omega_{i}\cap B_{\varepsilon}(\bar{x})$ and $A_{i}\in\Xi_{i}$ $(i=1,\ldots,n)$ such that $\max_{1\leq i\leq n}d(x_{i},A_{i})<\varepsilon\rho$ and $\bigcap_{i=1}^{n}(A_{i}-x_{i})\cap(\rho\mathbb{B})=\emptyset$.
  3. iii.
    The collection $\{\Xi_{1},\ldots,\Xi_{n}\}$ is stationary at $\bar{x}$ if it is approximately stationary at $\bar{x}$ with respect to $\{\{\bar{x}\},\ldots,\{\bar{x}\}\}$.
Theorem 1.2.
Let $\Xi_{1},\ldots,\Xi_{n}$ be families of closed subsets of a Banach space $X$, and $\bar{x}\in X$. Let $\Omega_{1},\ldots,\Omega_{n}\subset X$. If $\{\Xi_{1},\ldots,\Xi_{n}\}$ is approximately stationary at $\bar{x}$ with respect to $\{\Omega_{1},\ldots,\Omega_{n}\}$, then, for any $\varepsilon>0$ and $\tau\in(0,1)$, there exist $\bar{x}_{i}\in\Omega_{i}\cap B_{\varepsilon}(\bar{x})$, $A_{i}\in\Xi_{i}$, $x_{i}\in A_{i}\cap B_{\varepsilon}(\bar{x})$, $x_{i}^{*}\in N^{C}_{A_{i}}(x_{i})$ $(i=1,\ldots,n)$ and $x\in\varepsilon\mathbb{B}$, such that $\|\sum_{i=1}^{n}x_{i}^{*}\|<\varepsilon$, $\sum_{i=1}^{n}\|x_{i}^{*}\|=1$, and
$\sum_{i=1}^{n}\langle x_{i}^{*},x+\bar{x}_{i}-x_{i}\rangle>\tau\max_{1\leq i\leq n}\|x+\bar{x}_{i}-x_{i}\|.$ (1)
If $X$ is Asplund, then $N^{C}$ in the above assertion can be replaced by $N^{F}$.
The necessary conditions are formulated in Theorem 1.2 for the approximate stationarity, and, as a consequence, are also valid for the stationarity and extremality. Observe that, unlike the case of conventional linear perturbations, in the framework of general perturbations the fuzzy Fréchet separation property is necessary but not in general sufficient for approximate stationarity in Asplund spaces.
The main applications in [7, 8, 6, 9, 10, 11] of the set-valued version of the extremal principle were to multiobjective and set-valued optimization, multiplayer game, and optimal control problems. The more general (and simpler) model in Theorem 1.2 is capable of treating a wider range of such problems. In Sections 3 and 4, we demonstrate the applicability of Theorem 1.2 and its corollaries to set-valued optimization problems with general preference relations. This allows us to expand the range of set-valued optimization models studied in earlier publications, weaken their assumptions and streamline the proofs.
The structure of the paper is as follows. Section 2 recalls some definitions and facts used throughout the paper, particularly, the Ekeland variational principle and two types of subdifferential sum rules. The applicability of Theorem 1.2 is illustrated in Sections 3 and 4 considering set-valued optimization problems with general preference relations. A model with a single set-valued mapping is studied in detail in Section 3. Dual necessary conditions for approximate stationarity, and hence, also extremality, are proved in the form of a kind of multiplier rule under certain natural qualification conditions. A more general model with multiple set-valued mappings is briefly discussed in Section 4.
2 Preliminaries
[…]
The interior and closure of a set $\Omega$ are denoted by ${\rm int}\,\Omega$ and ${\rm cl}\,\Omega$, respectively. The distance from a point $x\in X$ to a subset $\Omega\subset X$ is defined by $d(x,\Omega):=\inf_{u\in\Omega}\|u-x\|$, and we use the convention $d(x,\emptyset)=+\infty$. Given a subset $\Omega$ of a normed space $X$ and a point $\bar{x}\in\Omega$, the sets (cf. [16, 17])
$N_{\Omega}^{F}(\bar{x}):=\Big{\{}x^{\ast}\in X^{\ast}\mid\limsup_{\Omega\ni x{\rightarrow}\bar{x},\;x\neq\bar{x}}\frac{\langle x^{\ast},x-\bar{x}}{\|x-\bar{x}\|}\leq 0\Big{\}},$ (2)
$N_{\Omega}^{C}(\bar{x}):=\left\{x^{\ast}\in X^{\ast}\mid\left\langle x^{\ast},z\right\rangle\leq 0\quad\mbox{for all}\quad z\in T_{\Omega}^{C}(\bar{x})\right\}$ (3)
are the Fréchet and Clarke normal cones to $\Omega$ at $\bar{x}$, where $T_{\Omega}^{C}(\bar{x})$ stands for the Clarke tangent cone to $\Omega$ at $\bar{x}$:
$T_{\Omega}^{C}(\bar{x}):=\big{\{}z\in X\mid\forall x_{k}{\rightarrow}\bar{x},\;x_{k}\in\Omega,\;\forall t_{k}\downarrow 0,\;\exists z_{k}\to z\\ \mbox{such that}\quad x_{k}+t_{k}z_{k}\in\Omega\quad\mbox{for all}\quad k\in\mathbb{N}\big{\}}.$
The sets (2) and (3) are nonempty closed convex cones satisfying $N_{\Omega}^{F}(\bar{x})\subset N_{\Omega}^{C}(\bar{x})$. If $\Omega$ is a convex set, they reduce to the normal cone in the sense of convex analysis:
$N_{\Omega}(\bar{x}):=\left\{x^{*}\in X^{*}\mid\langle x^{*},x-\bar{x}\rangle\leq 0\quad\mbox{for all}\quad x\in\Omega\right\}.$
For an extended-real-valued function $f:X\to\mathbb{R}_{\infty}:=\mathbb{R}\cup\{+\infty\}$ on a normed space $X$, its domain and epigraph are defined, respectively, by ${\rm dom}\,f:=\{x\in X\mid f(x)<+\infty\}$ and ${\rm epi}\,f:=\{(x,\alpha)\in X\times\mathbb{R}\mid f(x)\leq\alpha\}$. The Fréchet and Clarke subdifferentials of $f$ at $\bar{x}$ are defined by
$\partial^{F}f(\bar{x})=\left\{x^{*}\in X^{*}\mid(x^{*},-1)\in N^{F}_{{\rm epi}\,f}(\bar{x},f(\bar{x}))\right\},$ (4)
$\partial^{C}{f}(\bar{x})=\left\{x^{\ast}\in X^{\ast}\mid(x^{*},-1)\in N_{{\rm epi}\,f}^{C}(\bar{x},f(\bar{x}))\right\}.$ (5)
The sets (4) and (5) are closed and convex, and satisfy $\partial^{F}{f}(\bar{x})\subset\partial^{C}{f}(\bar{x})$. If $f$ is convex, they reduce to the subdifferential in the sense of convex analysis. In this case, we simply write $\partial{f}(\bar{x})$.
By convention, we set $N_{\Omega}^{F}(\bar{x})=N_{\Omega}^{C}(\bar{x}):=\emptyset$ if $\bar{x}\notin\Omega$ and $\partial^{F}{f}(\bar{x})=\partial^{C}{f}(\bar{x}):=\emptyset$ if $\bar{x}\notin{\rm dom}\,f$. It is easy to check that $N_{\Omega}^{F}(\bar{x})=\partial^{F}i_{\Omega}(\bar{x})$ and $N_{\Omega}^{C}(\bar{x})=\partial^{C}i_{\Omega}(\bar{x})$, where $i_{\Omega}$ is the indicator function of $\Omega$: $i_{\Omega}(x)=0$ if $x\in\Omega$ and $i_{\Omega}(x)=+\infty$ if $x\notin\Omega$.
A set-valued mapping $F:X\rightrightarrows Y$ between two sets $X$ and $Y$ is a mapping, which assigns to every $x\in X$ a (possibly empty) subset $F(x)$ of $Y$. We use the notations ${\rm gph}\,F:=\{(x,y)\in X\times Y\mid y\in F(x)\}$ and ${\rm dom}\,\>F:=\{x\in X\mid F(x)\neq\emptyset\}$ for the graph and the domain of $F$, respectively, and $F^{-1}:Y\rightrightarrows X$ for the inverse of $F$. This inverse (which always exists with possibly empty values at some $y$) is defined by $F^{-1}(y):=\{x\in X\mid y\in F(x)\}$, $y\in Y$. Obviously ${\rm dom}\,F^{-1}=F(X)$.
If $X$ and $Y$ are normed spaces, the Clarke coderivative $D^{*C}F(x,y)$ of $F$ at $(x,y)\in{\rm gph}\,F$ is a set-valued mapping defined by
$D^{*C}F(x,y)(y^{*}):=\{x^{*}\in X^{*}\mid(x^{*},-y^{*})\in N^{C}_{{\rm gph}\,F}(x,y)\},\quad y^{*}\in Y^{*}.$ (6)
Replacing the Clarke normal cone in (6) by the Fréchet one, we obtain the definition of the Fréchet coderivative.
Definition 2.1 (Aubin property).
A mapping $F:X\rightrightarrows Y$ between metric spaces has the Aubin property at $(\bar{x},\bar{y})\in{\rm gph}\,F$ if there exist $\tau>0$ and $\delta>0$ such that
$d(y,F(x))\leq\tau d(x,x^{\prime})\quad\mbox{for all}\quad x,x^{\prime}\in B_{\delta}(\bar{x}),\;y\in F(x^{\prime})\cap B_{\delta}(\bar{y}).$ (7)
The number $\tau$ is called the modulus of the Aubin property at $(\bar{x},\bar{y})$.
Lemma 2.2.
Let $X$ and $Y$ be normed spaces, $F:X\rightrightarrows Y$, and $(\bar{x},\bar{y})\in{\rm gph}\,F$.
  1. i.
    If $F$ has the Aubin property at $(\bar{x},\bar{y})$ with modulus $\tau>0$, then there is a $\delta>0$ such that
    $\|x^{*}\|\leq\tau\|y^{*}\|\quad\mbox{for all}\quad(x,y)\in{\rm gph}\,F\cap B_{\delta}(\bar{x},\bar{y}),\;(x^{*},y^{*})\in N^{F}_{{\rm gph}\,F}(x,y).$ (8)
  2. ii.
    If $\dim Y<+\infty$, then $N^{F}$ in the above assertion can be replaced by $N^{C}$.
3 Set-valued optimization
Let $X$ and $Y$ be normed spaces, $\Omega\subset X$, $K\subset Y$, $F:X\rightrightarrows Y$, $\bar{x}\in\Omega$ and $\bar{y}\in F(\bar{x})\cap K$. The set $K$ can be interpreted as the set of points in $Y$ “not greater” than $\bar{y}$ in the sense of some preference relation on $Y$ (as illustrated by Example 3.1). The triple $\{F,\Omega,K\}$ can model various set-valued optimization problems or constraint systems, with the mapping $F$ corresponding to the objective and/or constraint mappings. In this section, we discuss extremality and stationarity properties of $\{F,\Omega,K\}$.
Define two sets in the product space $X\times Y$:
[…]
Thus, $(\bar{x},\bar{y})\in\Omega_{1}\cap\Omega_{2}$, and we are in the setting of Theorem 1.2 (with $n=2$).
Example 3.1.
Let $Y:=\mathbb{R}$, $F(x):=\{\mu\in\mathbb{R}\mid f(x)\leq\mu\}$ for all $x\in X$, where $f:X\to\mathbb{R}_{\infty}$, $\bar{x}\in{\rm dom}\,f$, $\bar{y}:=f(\bar{x})$, and $K:=(-\infty,\bar{y}]$. If $\bar{x}$ is a local minimum of $f$ on $\Omega\subset X$, then $\{\Omega_{1},\Omega_{2}\}$ is extremal at $(\bar{x},\bar{y})$ in the sense of Definition ….
Our aim is to analyze extremality and stationarity properties of the pair of sets (9). When defining perturbations of $\Omega_{1}$ and $\Omega_{2}$, we are going to perturb only the $K$ component of $\Omega_{2}$. Let $\Xi$ denote a nonempty family of subsets of $Y$. We define two families of subsets of $X\times Y$: $\Xi_{1}:=\{\Omega_{1}\}$ (a single set) and $\Xi_{2}:=\{\Omega\times\widetilde{K}\mid\widetilde{K}\in\Xi\}$. Given a $y\in K$ and an $\varepsilon>0$, we denote $\Xi_{\varepsilon}(y):=\{\widetilde{K}\in\Xi\mid d(y,\widetilde{K})<\varepsilon\}$.
The properties in the next definition are reformulations of the corresponding ones in Definition 1.1 in the setting discussed above.
Definition 3.2.
The triple $\{F,\Omega,K\}$ is
  1. i.
    extremal at $(\bar{x},\bar{y})$ with respect to $\Xi$ if there is a $\rho\in(0,+\infty]$ such that, for any $\varepsilon>0$, there exists a $\widetilde{K}\in\Xi_{\varepsilon}(\bar{y})$ such that $F(\Omega\cap B_{\rho}(\bar{x}))\cap\widetilde{K}\cap B_{\rho}(\bar{y})=\emptyset$;
  2. ii.
    approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$ if, for any $\varepsilon>0$, there exist $\rho\in(0,\varepsilon)$, $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $y_{2}\in K\cap B_{\varepsilon}(\bar{y})$, and a $\widetilde{K}\in\Xi_{\varepsilon\rho}(y_{2})$ such that
    $F(x_{1}+(\Omega-x_{2})\cap(\rho\mathbb{B}_{X}))\cap(y_{1}+(\widetilde{K}-y_{2})\cap(\rho\mathbb{B}_{Y}))=\emptyset.$
Definition 3.2 (i) covers both local ($\rho<+\infty$) and global ($\rho=+\infty$) extremality. In the particular case of the property in part (ii) of Definition 3.2 when $x_{1}=x_{2}=\bar{x}$ and $y_{1}=y_{2}=\bar{y}$, the last equality in the definition takes the form $F(\Omega\cap B_{\rho}(\bar{x}))\cap\widetilde{K}\cap B_{\rho}(\bar{y})=\emptyset$, and we say that $\{F,\Omega,K\}$ is stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$. The next two examples illustrate the relationship between the properties in Definition 3.2.
Example 3.3.
Let $X=Y:=\mathbb{R}$, $\Omega:=[-2,2]$, $K:=(-\infty,0]$, $\Xi:=\{(-\infty,t]\mid t\in\mathbb{R}\}$, and $F_{1},F_{2},F_{3},F_{4}:\mathbb{R}\rightrightarrows\mathbb{R}$ be given by
$F_{1}(x)$ $:=[0,+\infty)\quad\mbox{for all}\;x\in\mathbb{R},\quad\;F_{2}(x):=\begin{cases}[x+1,+\infty)&\mbox{if }x<-1,\\ [0,+\infty)&\mbox{if }x\geq-1,\end{cases}$
$F_{3}(x)$ $:=[-x^{2},+\infty)\;\mbox{for all}\;x\in\mathbb{R},\quad F_{4}(x):=\begin{cases}[x,+\infty)&\mbox{if }x<0,\\ [-x^{2},+\infty)&\mbox{if }x\geq 0.\end{cases}$
Then $0\in\Omega$, $0\in F_{i}(0)$ for all $i=1,2,3,4$, and the following assertions hold true:
  1. i.
    $\{F_{1},\Omega,K\}$ is extremal at $(0,0)$ with respect to $\Xi$ with $\rho=+\infty$;
  2. ii.
    $\{F_{2},\Omega,K\}$ is locally extremal, but not extremal at $(0,0)$ with respect to $\Xi$ with $\rho=+\infty$;
  3. iii.
    $\{F_{3},\Omega,K\}$ is stationary, but not locally extremal at $(0,0)$ with respect to $\Xi$;
  4. iv.
    $\{F_{4},\Omega,K\}$ is approximately stationary, but not stationary at $(0,0)$ with respect to $\Xi$.
The assertions are straightforward. We only prove assertion (iv). Let $\varepsilon\in(0,1)$. Choose any $\rho\in(0,\varepsilon)$ and $t\in(\varepsilon\rho,\rho)$. Then $-t\in\rho\mathbb{B}_{\mathbb{R}}\subset\Omega$, $-t\in F_{4}(-t)$ and $-t\in\widetilde{K}\cap(\rho\mathbb{B}_{\mathbb{R}})$ for any $\widetilde{K}\in\Xi_{\varepsilon\rho}(0)$. Hence, $\{F_{4},\Omega,K\}$ is not stationary at $(0,0)$ with respect to $\Xi$.
Let $\varepsilon>0$. Choose a $\rho\in(0,\min\{\varepsilon,1\}/3)$ and points $(x_{1},y_{1}):=(\rho,-\rho^{2})\in{\rm gph}\,F_{4}\cap(\varepsilon\mathbb{B}_{\mathbb{R}^{2}})$, $x_{2}:=0\in\Omega\cap(\varepsilon\mathbb{B}_{\mathbb{R}})$, $y_{2}:=0\in K\cap(\varepsilon\mathbb{B}_{\mathbb{R}})$. Set $\widetilde{K}:=(-\infty,-3\rho^{2}]$ and observe that $\widetilde{K}\in\Xi_{\varepsilon\rho}(0)$. Then
$F_{4}(x_{1}+(\Omega-x_{2})\cap(\rho\mathbb{B}_{\mathbb{R}}))=F_{4}(0,2\rho)=(-4\rho^{2},+\infty),$
$y_{1}+(\widetilde{K}-y_{2})\cap(\rho\mathbb{B}_{\mathbb{R}})=-\rho^{2}+(-\rho,-3\rho^{2}]=(-\rho^{2}-\rho,-4\rho^{2}],$
and consequently, $F_{4}(x_{1}+(\Omega-x_{2})\cap(\rho\mathbb{B}_{\mathbb{R}}))\cap(y_{1}+(\widetilde{K}-y_{2})\cap(\rho\mathbb{B}_{\mathbb{R}}))=\emptyset$. Thus, $\{F_{4},\Omega,K\}$ is approximately stationary at $(0,0)$ with respect to $\Xi$.
Example 3.4.
Let $X=Y:=\mathbb{R}$, $\Omega:=\mathbb{R}$, $K:=(-\infty,0]$ and $F:\mathbb{R}\rightrightarrows\mathbb{R}$ be given by
$F(x):=\begin{cases}\{x\sqrt{2}\}&\mbox{if }x\text{ is rational},\\ \{x\}&\mbox{otherwise}.\end{cases}$
Then $0\in F(0)$. The triple $\{F,\Omega,K\}$ is not approximately stationary at $(0,0)$ with respect to $\Xi:=\{(-\infty,t]\mid t\in\mathbb{R}\}$. Indeed, let $\varepsilon\in(0,1)$. Choose any $\rho\in(0,\varepsilon)$, $(x_{1},y_{1})\in{\rm gph}\,F$, $x_{2}\in\mathbb{R}$, $y_{2}\in(-\varepsilon,0]$ and $\widetilde{K}\in\Xi_{\varepsilon\rho}(y_{2})$. Then $x_{1}+(\Omega-x_{2})\cap(\rho\mathbb{B}_{X})=B_{\rho}(x_{1})$, $\widetilde{K}=(-\infty,t]$ for some $t>y_{2}-\varepsilon\rho$ and $y_{1}+(\widetilde{K}-y_{2})\cap(\rho\mathbb{B}_{Y})\supset(y_{1}-\rho,y_{1}-\varepsilon\rho)$. If $x_{1}$ is rational, then $y_{1}=x_{1}\sqrt{2}$, and choosing a rational number $\hat{x}\in(x_{1}-\rho/\sqrt{2},x_{1}-\varepsilon\rho/\sqrt{2})\subset B_{\rho}(x_{1})$, we get $\hat{y}:=\hat{x}\sqrt{2}\in F(\hat{x})\cap(y_{1}-\rho,y_{1}-\varepsilon\rho)$. If $x_{1}$ is irrational, then $y_{1}=x_{1}$, and choosing an irrational number $\hat{x}\in(x_{1}-\rho,x_{1}-\varepsilon\rho)\subset B_{\rho}(x_{1})$, we get $\hat{y}:=\hat{x}\in F(\hat{x})\cap(y_{1}-\rho,y_{1}-\varepsilon\rho)$. Hence, $\{F,\Omega,K\}$ is not approximately stationary at $(0,0)$ with respect to $\Xi$. At the same time, it is extremal (with $\rho=+\infty$) with respect to $\Xi:=\{\{-1/n\}\mid n\in\mathbb{N}\}$ because $F(X)\cap\widetilde{K}=\emptyset$ for all $\widetilde{K}\in\Xi$.
Application of Theorem 1.2 yields necessary conditions for approximate stationarity, and hence, also extremality.
Theorem 3.5.
Let $X$ and $Y$ be Banach spaces, the sets $\Omega$, ${\rm gph}\,F$ and all members of $\Xi$ be closed. If the triple $\{F,\Omega,K\}$ is approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$, then, for any $\varepsilon>0$, there exist $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $\widetilde{K}\in\Xi$, $y_{2}\in\widetilde{K}\cap B_{\varepsilon}(\bar{y})$, $(x_{1}^{*},y_{1}^{*})\in N^{C}_{{\rm gph}\,F}(x_{1},y_{1})$, $x_{2}^{*}\in N^{C}_{\Omega}(x_{2})$ and $y_{2}^{*}\in N^{C}_{\widetilde{K}}(y_{2})$ such that $\|(x_{1}^{*},y_{1}^{*})+(x_{2}^{*},y_{2}^{*})\|<\varepsilon$ and $\|(x_{1}^{*},y_{1}^{*})\|+\|(x_{2}^{*},y_{2}^{*})\|=1$.
If $X$ is Asplund, then $N^{C}$ in the above assertion can be replaced by $N^{F}$.
The normalization condition $\|(x_{1}^{*},y_{1}^{*})\|+\|(x_{2}^{*},y_{2}^{*})\|=1$ in Theorem 3.5 ensures that normal vectors $(x_{1}^{*},y_{1}^{*})$ to ${\rm gph}\,F$ remain sufficiently large when $\varepsilon\downarrow 0$, i.e., $x_{1}^{*}$ and $y_{1}^{*}$ cannot go to 00 simultaneously. The case when vector $y_{1}^{*}$ are bounded away from 00 (hence, one can assume $\|y_{1}^{*}\|=1$) is of special interest as it leads to a proper multiplier rule. A closer look at the alternative: either $y_{1}^{*}$ is bounded away from 00 as $\varepsilon\downarrow 0$, or it is not (hence, $x_{1}^{*}$ remains large), allows one to formulate the following consequence of Theorem 3.5.
Corollary 3.6.
Let $X$ and $Y$ be Banach spaces, the sets $\Omega$, ${\rm gph}\,F$ and all members of $\Xi$ be closed. If the triple $\{F,\Omega,K\}$ is approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$, then one of the following assertions holds true:
  1. i.
    there is an $M>0$ such that, for any $\varepsilon>0$, there exist $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $\widetilde{K}\in\Xi$, $y_{2}\in\widetilde{K}\cap B_{\varepsilon}(\bar{y})$, and $y^{*}\in N^{C}_{\widetilde{K}}(y_{2})+\varepsilon\mathbb{B}_{Y^{*}}$ such that $\|y^{*}\|=1$ and
    $0\in D^{*C}F(x_{1},y_{1})(y^{*})+N^{C}_{\Omega}(x_{2})\cap(M\mathbb{B}_{X^{*}})+\varepsilon\mathbb{B}_{X^{*}};$ (10)
  2. ii.
    for any $\varepsilon>0$, there exist $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $x_{1}^{*}\in D^{*C}F(x_{1},y_{1})(\varepsilon\mathbb{B}_{Y^{*}})$ and $x_{2}^{*}\in N^{C}_{\Omega}(x_{2})$ such that $\|x_{1}^{*}\|+\|x_{2}^{*}\|=1$ and $\|x_{1}^{*}+x_{2}^{*}\|<\varepsilon$.
If $X$ is Asplund, then $N^{C}$ and $D^{*C}$ in the above assertions can be replaced by $N^{F}$ and $D^{*F}$, respectively.
Proof.
Let the triple $\{F,\Omega,K\}$ be approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$. By Theorem 3.5, for any $j\in\mathbb{N}$, there exist $(x_{1j},y_{1j})\in{\rm gph}\,F\cap B_{1/j}(\bar{x},\bar{y})$, $x_{2j}\in\Omega\cap B_{1/j}(\bar{x})$, ${\widetilde{K}_{j}}\in\Xi$, $y_{2j}\in\widetilde{K}_{j}\cap B_{1/j}(\bar{y})$, $(x_{1j}^{*},y_{1j}^{*})\in N^{C}_{{\rm gph}\,F}(x_{1j},y_{1j})$, $x_{2j}^{*}\in N^{C}_{\Omega}(x_{2j})$ and $y^{*}_{2j}\in N^{C}_{\widetilde{K}_{j}}(y_{2j})$ such that $\|(x_{1j}^{*},y_{1j}^{*})\|+\|(x_{2j}^{*},y_{2j}^{*})\|=1$ and $\|(x_{1j}^{*},y_{1j}^{*})+(x_{2j}^{*},y_{2j}^{*})\|<1/j$. We consider two cases.
Case 1. $\limsup_{j\to+\infty}\|y_{1j}^{*}\|>\alpha>0$. Note that $\alpha<1$. Set $M:=1/\alpha$. Let $\varepsilon>0$. Choose a number $j\in\mathbb{N}$ so that $j^{-1}<\alpha\varepsilon$ and $\|y_{1j}^{*}\|>\alpha$. Set $x_{1}:=x_{1j}$, $y_{1}:=y_{1j}$, $x_{2}:=x_{2j}$, $\widetilde{K}:=\widetilde{K}_{j}$, $y_{2}:=y_{2j}$, $y^{*}:=-y_{1j}^{*}/\|y_{1j}^{*}\|$, $x_{1}^{*}:=x_{1j}^{*}/\|y_{1j}^{*}\|$, $x_{2}^{*}:=x_{2j}^{*}/\|y_{1j}^{*}\|$ and $y_{2}^{*}:=y_{2j}^{*}/\|y_{1j}^{*}\|$. Then $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $y_{2}\in\widetilde{K}\cap B_{\varepsilon}(\bar{y})$, $x_{1}^{*}\in D^{*C}F(x_{1},y_{1})(y^{*})$, $x_{2}^{*}\in N^{C}_{\Omega}(x_{2})$, ${\|y^{*}\|=1}$, $y^{*}_{2}\in N^{C}_{\widetilde{K}}(y_{2})$, $\|x_{2}^{*}\|<1/\alpha=M$. Furthermore, $\|y^{*}-y_{2}^{*}\|=\|y_{1j}^{*}+y_{2j}^{*}\|/\|y_{1j}^{*}\|<1/(\alpha j)<\varepsilon$, hence, $y^{*}\in N^{C}_{\widetilde{K}}(y_{2})+\varepsilon\mathbb{B}_{Y^{*}}$; and $\|x_{1}^{*}+x_{2}^{*}\|=\|x_{1j}^{*}+x_{2j}^{*}\|/\|y_{1j}^{*}\|<1/(\alpha j)<\varepsilon$, hence, condition (10) is satisfied. Thus, assertion (i) holds true.
Case 2. $\lim_{j\to+\infty}\|y_{1j}^{*}\|=0$. Then $y_{2j}^{*}\to 0$, $x_{1j}^{*}+x_{2j}^{*}\to 0$ and $1\geq\|x_{1j}^{*}\|+\|x_{2j}^{*}\|\to 1$ as $j\to+\infty$. Let $\varepsilon>0$. Choose a number $j\in\mathbb{N}$ so that $\|x_{1j}^{*}\|+\|x_{2j}^{*}\|>0$ and $\max\{j^{-1},\|y_{1j}^{*}\|,\|x_{1j}^{*}+x_{2j}^{*}\|/(\|x_{1j}^{*}\|+\|x_{2j}^{*}\|)\}<\varepsilon$. Set $x_{1}:=x_{1j}$, $y_{1}:=y_{1j}$, $x_{2}:=x_{2j}$, $x_{1}^{*}:=x_{1j}^{*}/(\|x_{1j}^{*}\|+\|x_{2j}^{*}\|)$ and $x_{2}^{*}:=x_{2j}^{*}/(\|x_{1j}^{*}\|+\|x_{2j}^{*}\|)$. Then $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $x_{1}^{*}\in D^{*C}F(x_{1},y_{1})(\varepsilon\mathbb{B}_{Y^{*}})$, $x_{2}^{*}\in N^{C}_{\Omega}(x_{2})$, $\|x_{1}^{*}\|+\|x_{2}^{*}\|=1$, and $\|x_{1}^{*}+x_{2}^{*}\|<\varepsilon$. Thus, assertion (ii) holds true.
If $X$ is Asplund, then $N^{C}$ and $D^{*C}$ in the above arguments can be replaced by $N^{F}$ and $D^{*F}$, respectively. ∎
Remark 1.
  1. i.
    Part (i) of Corollary 3.6 gives a kind of multiplier rule with $y^{*}$ playing the role of the vector of multipliers.
  2. ii.
    Part (ii) corresponds to ‘singular’ behavior of $F$ on $\Omega$. It involves ‘horizontal’ normals to the graph of $F$; the $y^{*}$ component vanishes, and consequently, $K$ and $\widetilde{K}$ play no role.
The following condition is the negation of the condition in Corollary 3.6 (ii).
    1. $(QC)_{C}$
      there is an $\varepsilon>0$ such that $\|x_{1}^{*}+x_{2}^{*}\|\geq\varepsilon$ for all $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $x_{1}^{*}\in D^{*C}F(x_{1},y_{1})(\varepsilon\mathbb{B}_{Y^{*}})$ and $x_{2}^{*}\in N^{C}_{\Omega}(x_{2}))$ such that $\|x_{1}^{*}\|+\|x_{2}^{*}\|=1$.
It excludes the singular behavior mentioned in Remark 1 (ii) and serves as a qualification condition ensuring that only the condition in part (i) is possible. We denote by $(QC)_{F}$ the analogue of $(QC)_{C}$ with $N^{F}$ and $D^{*F}$ in place of $N^{C}$ and $D^{*C}$, respectively.
Corollary 3.7.
Let $X$ and $Y$ be Banach spaces, $\Omega$, ${\rm gph}\,F$ and all members of $\Xi$ be closed. Suppose that the triple $\{F,\Omega,K\}$ is approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$. If condition $(QC)_{C}$ is satisfied, then assertion (i) in Corollary 3.6 holds true.
If $X$ is Asplund and condition $(QC)_{F}$ is satisfied, then assertion (i) in Corollary 3.6 holds true with $N^{F}$ and $D^{*F}$ in place of $N^{C}$ and $D^{*C}$, respectively.
The next proposition provides two typical sufficient conditions for the fulfillment of conditions $(QC)_{C}$ and $(QC)_{F}$.
Proposition 3.8.
Let $X$ and $Y$ be normed spaces.
  1. i.
    If $F$ has the Aubin property at $(\bar{x},\bar{y})$, then $(QC)_{F}$ is satisfied. If, additionally, $\dim Y<+\infty$, then $(QC)_{C}$ is satisfied too.
  2. ii.
    If $\bar{x}\in{\rm int}\,\Omega$, then both $(QC)_{C}$ and $(QC)_{F}$ are satisfied.
Proof.
  1. i.
    If $F$ has the Aubin property at $(\bar{x},\bar{y})$, then, by Lemma 2.2, condition (8) is satisfied with some $\tau>0$ and $\delta>0$, and, if $\dim Y<+\infty$, then the latter condition is also satisfied with $N^{C}$ in place of $N^{F}$. Hence, $(QC)_{F}$ is satisfied with $\varepsilon:=1/(2\tau+1)$, as well as $(QC)_{C}$ if $\dim Y<+\infty$. Indeed, if $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\delta}(\bar{x},\bar{y})$ and $x_{1}^{*}\in D^{*}F(x_{1},y_{1})(y_{1}^{*})$ (where $D^{*}$ stands for either $D^{*C}$ or $D^{*F}$), $x_{2}^{*}\in X^{*}$, $\|x_{1}^{*}\|+\|x_{2}^{*}\|=1$ and $\|y_{1}^{*}\|<\varepsilon$, then $\|x_{1}^{*}+x_{2}^{*}\|\geq\|x_{2}^{*}\|-\|x_{1}^{*}\|=1-2\|x_{1}^{*}\|>1-2\tau\varepsilon=\varepsilon.$
  2. ii.
    If $\bar{x}\in{\rm int}\,\Omega$, then $N^{C}_{\Omega}(x_{2})=N^{F}_{\Omega}(x_{2})=\{0\}$ for all $x_{2}$ near $\bar{x}$, and consequently, for any normal vector $x_{2}^{*}$ to $\Omega$ at $x_{2}$ and any $x_{1}^{*}\in X^{*}$, condition $\|x_{1}^{*}\|+\|x_{2}^{*}\|=1$ yields $\|x_{1}^{*}+x_{2}^{*}\|=1$. Hence, both $(QC)_{C}$ and $(QC)_{F}$ are satisfied with any sufficiently small $\varepsilon$.
Remark 2.
In view of Proposition 3.8, each of the Corollaries 3.6 and 3.7 covers [18, Theorem 3.1] and [9, Corollary 3.1], which give dual necessary conditions for Pareto optimality, as well as [8, Proposition 5.1] for a vector optimization problem with a general preference relation.
The next example illustrates the verification of the necessary conditions for approximate stationarity in Corollaries 3.6 and 3.7 for the triple $\{F_{4},\Omega,K\}$ in Example 3.3.
Example 3.9.
Let $X=Y:=\mathbb{R}$, and $\Omega$, $K$, $\Xi$ and $F_{4}$ be as in Example 3.3. Thus, the triple $\{F_{4},\Omega,K\}$ is approximately stationary at $(0,0)$ with respect to $\Xi$, and the conclusions of Corollary 3.6 must hold true. Moreover, the assumptions in both parts of Proposition 3.8 are satisfied, and consequently, condition $(QC)_{F}$ holds true. By Corollary 3.7, assertion (i) in Corollary 3.6 holds true with $N^{F}$ and $D^{*F}$ in place of $N^{C}$ and $D^{*C}$, respectively. We now verify this assertion.
Let $M>0$ and $\varepsilon>0$. Choose a $t\in(0,\min\{\varepsilon,1\})$. Set $(x_{1},y_{1}):=(t/2,-t^{2}/4)\in{\rm gph}\,F_{4}\cap(\varepsilon\mathbb{B}_{\mathbb{R}^{2}})$, $x_{2}:=0\in\Omega\cap{(\varepsilon\mathbb{B}_{\mathbb{R}})}$, $\widetilde{K}:=(-\infty,-t]\in\Xi$, $y_{2}:=-t\in\widetilde{K}\cap{(\varepsilon\mathbb{B}_{\mathbb{R}})}$ and $y^{*}:=1$. Thus, $N^{F}_{\Omega}(x_{2})=\{0\}$, $y^{*}\in N^{F}_{\widetilde{K}}(y_{2})$ and $D^{*F}F_{4}(x_{1},y_{1})(y^{*})=\{-t\}$. Hence,
$D^{*F}F_{4}(x_{1},y_{1})(y^{*})+N^{F}_{\Omega}(x_{2})\cap(M\mathbb{B}_{\mathbb{R}})=\{-t\}\in\varepsilon\mathbb{B}_{\mathbb{R}},$
i.e., assertion (i) in Corollary 3.6 holds true (in terms of Fréchet normals and coderivatives).
We now consider a particular case of the model (9) when the set $K$ and its perturbations are determined by an abstract level-set mapping $L:Y\rightrightarrows Y$. The latter mapping defines a preference relation $\prec$ on $Y$: $v\prec y$ if and only if $v\in L(y)$; see, e.g., [19, p. 67].
Given a point $y\in Y$ and a number $\varepsilon>0$, we employ below the following notations:
$L^{\circ}(y):=L(y)\setminus\{y\},$ (11)
$L^{\bullet}_{\varepsilon}(y):=L^{-1}(B_{\varepsilon}(y))=\{v\in Y\mid d(y,L(v))<\varepsilon\}.$ (12)
Certain requirements are usually imposed on $L$ in order to make the corresponding preference relation useful in optimization and applications; see, e.g., [7, 8, 6, 19]. In this section, we are going to make use of the following properties of $L$ at or near the reference point $\bar{y}$:
  1. O1.
    $\liminf\limits_{L^{\circ}(\bar{y})\ni y\to\bar{y}}d(\bar{y},L(y))=0$;
  2. O2.
    $\bar{y}\in{\rm cl}\,L^{\circ}(\bar{y})$;
  3. O3.
    $\bar{y}\notin L(\bar{y})$;
  4. O4.
    $y\in{\rm cl}\,L(y)$ for all $y$ near $\bar{y}$;
  5. O5.
    if $y\in L^{\circ}(\bar{y})$ and $v\in{\rm cl}\,L(y)$, then $v\in L^{\circ}(\bar{y})$;
  6. O6.
    if $y\in L(\bar{y})$ and $v\in{\rm cl}\,L(y)$, then $v\in L(\bar{y})$.
Some characterizations of the properties and relations between them are collected in the next proposition.
Proposition 3.10.
Let $L:Y\rightrightarrows Y$, and $L^{\circ}$ and $L^{\bullet}_{\varepsilon}$ be given by (11) and (12), respectively.
  1. i.
[…]
  3. ii.
[…]
  4. iii.
[…]
  5. iv.
[…]
  6. v.
[…]
  7. vi.
[…]
Proof.
  1. i.
[…]
  2. ii.
[…]
  3. iii.
    The assertion is a consequence of definition (11).
  4. iv.
    Suppose conditions (O2) and (O4) are satisfied. Let $\varepsilon>0$. Thanks to (O4), we can choose a $\xi\in(0,\varepsilon)$ such that $y\in{\rm cl}\,L(y)$ for all $y\in B_{\xi}(\bar{y})$. If $y\in B_{\xi}(\bar{y})$, then $d(\bar{y},L(y))=d(\bar{y},{\rm cl}\,L(y))\leq\|y-\bar{y}\|<\xi$, and consequently, $y\in L^{\bullet}_{\xi}(\bar{y})$. Thus, $B_{\xi}(\bar{y})\subset L^{\bullet}_{\xi}(\bar{y})$. Thanks to (O2), we have
    $L^{\bullet}_{\varepsilon}(\bar{y})\cap L^{\circ}(\bar{y})\cap B_{\varepsilon}(\bar{y})\supset L^{\bullet}_{\xi}(\bar{y})\cap L^{\circ}(\bar{y})\cap B_{\xi}(\bar{y})=L^{\circ}(\bar{y})\cap B_{\xi}(\bar{y})\neq\emptyset.$
    Since $\varepsilon$ is an arbitrary positive number, in view of (i), this proves (O1).
  5. v.
[…]
  6. vi.
    The assertion is a consequence of (iii).
Remark 3.
Properties (O4) and (O6) are the components of the definition of closed preference relation (see [6, Definition 5.55], [11, p. 583], [19, p. 68]) widely used in vector and set-valued optimization. They are called, respectively, local satiation (around $\bar{y}$) and almost transitivity. Note that the latter property is actually stronger than the conventional transitivity. It is not satisfied for the preference defined by the lexicographical order (see [6, Example 5.57]) and some other natural preference relations important in vector optimization and its applications including those to welfare economics (see [19, Sect. 15.3]). Closed preference relations are additionally assumed in [6, 11, 19] to be nonreflexive, thus, satisfying, in particular, property (O3). In view of Proposition 3.10, if a preference relation satisfies properties (O3), (O4) and (O6), it also satisfies properties (O1), (O2) and (O5). In this section, we employ the weaker properties (O1) and (O5), which are satisfied by most conventional and many unconventional preference relations. This makes our model applicable to a wider range of multiobjective and set-valued optimization problems compared to those studied in [6, 11, 19].
Example 3.11.
Let $L(y):=\{y\}$ for all $y\in Y$. Then $L^{\circ}(y)=\emptyset$ and $L_{\varepsilon}^{\bullet}(y)=B_{\varepsilon}(y)$ for all $y\in Y$ and $\varepsilon>0$. Thus, properties (O4) and (O5) are obviously satisfied, while properties (O1) and (O2) are violated.
Example 3.12.
Let $L:\mathbb{R}^{2}\rightrightarrows\mathbb{R}^{2}$ be defined by
$L(y_{1},y_{2}):=\begin{cases}\{(v_{1},v_{2})\in\mathbb{R}^{2}\mid v_{1}<y_{1},\;v_{2}<y_{2}\}&\text{if }(y_{1},y_{2})\neq(0,0),\\ \{(0,0)\}&\text{otherwise}.\end{cases}$
Then $L^{\circ}(y_{1},y_{2})=L(y_{1},y_{2})$ if $(y_{1},y_{2})\neq(0,0)$, $L^{\circ}(0,0)=\emptyset$ and $L_{\varepsilon}^{\bullet}(y_{1},y_{2})=\{(v_{1},v_{2})\in\mathbb{R}^{2}\mid\min\{y_{1}-v_{1},y_{2}-v_{2}\}<\varepsilon\}$ for all $(y_{1},y_{2})\in\mathbb{R}^{2}$ and $\varepsilon>0$. Let $\bar{y}:=(0,0)$. As in Example 3.11, properties (O4) and (O5) are satisfied, while properties (O1) and (O2) are violated.
We are going to employ in our model the set $K$ and, given a $\delta>0$, the family $\Xi^{\delta}$ of its perturbations defined as follows:
[…]
Then $\Xi^{\delta}_{\varepsilon}(y)=\{{\rm cl}\,L(v)\mid v\in L^{\bullet}_{\varepsilon}(y)\cap K\cap B_{\delta}(\bar{y})\}$ for all $y\in Y$, $\varepsilon>0$ and $\delta>0$. Note that members of $\Xi^{\delta}_{\varepsilon}$ are not simply translations (deformations) of the fixed set $L(\bar{y})$ (or $L^{\circ}(\bar{y})$); they are defined by sets $L(y)$ where $y$ does not have to be equal to $\bar{y}$.
Remark 4.
The set $K$ playing a crucial role in what follows is defined in (13) via the set $L(\bar{y})$. Conversely, given a set $K$ containing $\bar{y}$, one can naturally define the level-set mapping by $L(y)=y-\bar{y}+K$ for all $y\in Y$. Then the first definition in (13) reproduces the set $K$, while the second one defines the family of perturbations as the traditional collection of translations of ${\rm cl}\,K$, i.e., $\Xi^{\delta}=\{y-\bar{y}+{\rm cl}\,K\mid y\in K\cap B_{\delta}(\bar{y})\}$. In the rest of this section, we consider the more general model based on an arbitrary level-set mapping $L$.
The properties in Definition 3.2 take the following form.
Proposition 3.13.
Let $\delta>0$, and $K$ and $\Xi^{\delta}$ be given by (13). The triple $\{F,\Omega,K\}$ is
  1. i.
    extremal at $(\bar{x},\bar{y})$ with respect to $\Xi^{\delta}$ if and only if there is a $\rho>0$ such that, for any ${\varepsilon>0}$, there exists a $y\in L^{\bullet}_{\varepsilon}(\bar{y})\cap L^{\circ}(\bar{y})\cap B_{\delta}(\bar{y})$ such that $F(\Omega\cap B_{\rho}(\bar{x}))\cap{\rm cl}\,L(y)\cap B_{\rho}(\bar{y})=\emptyset$;
  2. ii.
    approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$ if and only if, for any $\varepsilon>0$, there exist a $\rho\in(0,\varepsilon)$ and points $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $y_{2}\in K\cap B_{\varepsilon}(\bar{y})$ and a $y\in L^{\bullet}_{\varepsilon\rho}(y_{2})\cap K\cap B_{\delta}(\bar{y})$ such that
    $F(x_{1}+(\Omega-x_{2})\cap(\rho\mathbb{B}_{X}))\cap(y_{1}+({\rm cl}\,L(y)-y_{2})\cap(\rho\mathbb{B}_{Y}))=\emptyset.$
The statements of Theorem 3.5 and its corollaries can be easily adjusted to the current setting. For instance, Corollary 3.7 can be reformulated as follows.
Corollary 3.14.
Let $X$ and $Y$ be Banach spaces, $\Omega$ and ${\rm gph}\,F$ be closed, $\delta>0$, and $K$ and $\Xi^{\delta}$ be given by (13). Suppose that the triple $\{F,\Omega,K\}$ is approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$. If condition $(QC)_{C}$ is satisfied, then there is an $M>0$ such that, for any $\varepsilon>0$, there exist $(x_{1},y_{1})\in{\rm gph}\,F\cap B_{\varepsilon}(\bar{x},\bar{y})$, $x_{2}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $y\in K\cap B_{\delta}(\bar{y})$, $y_{2}\in{\rm cl}\,L(y)\cap B_{\varepsilon}(\bar{y})$, and $y^{*}\in N^{C}_{{\rm cl}\,L(y)}(y_{2})+\varepsilon\mathbb{B}_{Y^{*}}$ such that $\|y^{*}\|=1$ and condition (10) holds true.
If $X$ is Asplund and condition $(QC)_{F}$ is satisfied, then the above assertion holds true with $N^{F}$ and $D^{*F}$ in place of $N^{C}$ and $D^{*C}$, respectively.
The properties in Definition 3.2 are rather general. They cover various optimality and stationarity concepts in vector and set-valued optimization.
With $\Omega$, $F$ and $L$ as above, and points $\bar{x}\in\Omega$ and $\bar{y}\in F(\bar{x})$, the next definition seems reasonable.
Definition 3.15.
The point $(\bar{x},\bar{y})$ is extremal for $F$ on $\Omega$ if there is a $\rho\in(0,+\infty]$ such that
$F(\Omega\cap B_{\rho}(\bar{x}))\cap L^{\circ}(\bar{y})\cap B_{\rho}(\bar{y})=\emptyset.$ (14)
Definition 3.15 covers both local ($\rho<+\infty$) and global ($\rho=+\infty$) extremality. The above concept is applicable, in particular, to solutions of the following set-valued minimization problem with respect to the preference determined by $L$:
minimize $\text{minimize }\;F(x)\quad\text{subject to }\;x\in\Omega.$ ($P$)
Note that the concept in Definition 3.15 is broader than just (local) minimality as $F$ is not assumed to be an objective mapping of an optimization problem. It can, for instance, be involved in modeling constraints.
Remark 5.
The property in Definition 3.15 is similar to the one in the definition of fully localized minimizer in [20, Definition 3.1] (see also [19, p. 68]). The latter definition uses the larger set ${\rm cl}\,L(\bar{y})\setminus\{\bar{y}\}$ (compare it with (11)) in place of $L^{\circ}(\bar{y})$ in (14). It is not difficult to check that the two properties are equivalent (when $\rho<+\infty$). Unlike many solution concepts in vector optimization, the above definition involves “image localization” (hence, is in general weaker). It has proved to be useful when studying local optimal allocations of welfare economics; cf. [20, 19].
We next show that, under some mild assumption on the level-set mapping $L$, the extremality in the sense of (14) can be treated in the framework of the extremality in the sense of Definition 3.2 (i) (or its characterization in Proposition 3.13 (i).
Proposition 3.16.
Let $\bar{x}\in\Omega$, $\bar{y}\in F(\bar{x})$, $\delta>0$, and $K$ and $\Xi^{\delta}$ be given by (13). Suppose $L$ satisfies conditions (O1) and (O5). If $(\bar{x},\bar{y})$ is extremal for $F$ on $\Omega$, then $\{F,\Omega,K\}$ is extremal at $(\bar{x},\bar{y})$ with respect to $\Xi^{\delta}$.
Proof.
In view of (O1), it follows from Proposition 3.10 (i) that
$L^{\bullet}_{\varepsilon}(\bar{y})\cap L^{\circ}(\bar{y})\cap B_{\delta}(\bar{y})\neq\emptyset\quad\mbox{for all}\quad\varepsilon>0\;\text{and}\;\delta>0.$ (15)
Suppose $\{F,\Omega,K\}$ is not extremal at $(\bar{x},\bar{y})$ with respect to $\Xi^{\delta}$ for some ${\delta>0}$. Let $\rho\in(0,+\infty]$. By Proposition 3.13 (i), there exists an $\varepsilon>0$ such that, for any $y\in L^{\bullet}_{\varepsilon}(\bar{y})\cap L^{\circ}(\bar{y})\cap B_{\delta}(\bar{y})$, it holds
$F(\Omega\cap B_{\rho}(\bar{x}))\cap{\rm cl}\,L(y)\cap B_{\rho}(\bar{y})\neq\emptyset.$ (16)
By (15), there is a point $y\in L^{\bullet}_{\varepsilon}(\bar{y})\cap L^{\circ}(\bar{y})\cap B_{\delta}(\bar{y})$, and we can choose a point $\hat{y}$ belonging to the set in (16). Thus, $y\in L^{\circ}(\bar{y})$ and $\hat{y}\in{\rm cl}\,L(y)$. Thanks to (O5), we have $\hat{y}\in L^{\circ}(\bar{y})$, and consequently, $\hat{y}\in F(\Omega\cap B_{\rho}(\bar{x}))\cap L^{\circ}(\bar{y})\cap B_{\rho}(\bar{y})$. Since $\rho\in(0,+\infty]$ is arbitrary, $(\bar{x},\bar{y})$ is not extremal for $F$ on $\Omega$. ∎
Thanks to Proposition 3.16, if the level-set mapping $L$ satisfies conditions (O1) and (O5), then extremal points of problem ($P$) satisfy the necessary conditions in Theorem 3.5 and its corollaries.
4 Set-valued optimization: multiple mappings
It is not difficult to upgrade the model used in Definition 3.2 and the subsequent statements to make it directly applicable to constraint optimization problems: instead of a single mapping $F:X\rightrightarrows Y$ and subset $K\subset Y$ with $\bar{y}\in F(\bar{x})\cap K$ for some $\bar{x}\in X$, one can consider a finite collection of mappings $F_{i}:X\rightrightarrows Y_{i}$ between normed spaces together with a collection of subsets $K_{i}\subset Y_{i}$ and points $\bar{y}_{i}\in F_{i}(\bar{x})\cap K_{i}$ $(i=1,\ldots,n)$.
This more general setting can be viewed as a structured particular case of the model $\{F,\Omega,K\}$ considered in Section 3 if one sets $Y:=Y_{1}\times\ldots\times Y_{n}$, $F:=(F_{1},\ldots,F_{n}):X\rightrightarrows Y$, $K:=K_{1}\times\ldots\times K_{n}$ and $\bar{y}:=(\bar{y}_{1},\ldots,\bar{y}_{n})$. To make full use of this structure, instead of the pair of sets (9) we are going to employ the family of $n+1$ subsets of $X\times Y_{1}\times\ldots\times Y_{n}$ defined as follows:
$\Omega_{i}:=\{(x,y_{1},\ldots,y_{n})\in X\times Y_{1}\times\ldots\times Y_{n}\mid y_{i}\in F_{i}(x)\}\quad(i=1,\ldots,n),$
$\Omega_{n+1}:=\Omega\times K.$
Observe that $(\bar{x},\bar{y})\in\bigcap_{i=1}^{n+1}\Omega_{i}$.
When defining perturbations of $\Omega_{1},\ldots,\Omega_{n+1}$, we are going to perturb only the $K$ component of $\Omega_{n+1}$. The family of perturbations of $K$ is defined as the product: $\Xi:=\Xi_{1}\times\ldots\times\Xi_{n}$, where, for each $i=1,\ldots,n$, $\Xi_{i}$ denotes a nonempty family of subsets of $Y_{i}$. The families of perturbations of $\Omega_{1},\ldots,\Omega_{n+1}$ are defined as follows: for each $i=1,\ldots,n$, we set $\widetilde{\Xi}_{i}:=\{\Omega_{i}\}$ (a single set), and $\widetilde{\Xi}_{n+1}:=\{\Omega\times\widetilde{K}\mid\widetilde{K}\in\Xi\}$. Given a $y\in K$ and an $\varepsilon>0$, we denote $\Xi_{\varepsilon}(y):=\{\widetilde{K}\in\Xi\mid d(y,\widetilde{K})<\varepsilon\}$.
With these notations, Definition 1.1, applied to the families $\Omega_{1},\ldots,\Omega_{n+1}$ and $\widetilde{\Xi}_{1},\ldots,\widetilde{\Xi}_{n+1}$, produces an extension of Definition 3.2. One can observe that in this more general setting part (i) of Definition 3.2 remains unchanged, while an extension of part (ii) takes the following form.
Definition 4.1.
The triple $\{F,\Omega,K\}$ defined above is approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$ if, for any $\varepsilon>0$, there exist a $\rho\in(0,\varepsilon)$, $(x_{i},y_{i})\in{\rm gph}\,F_{i}\cap B_{\varepsilon}(\bar{x},\bar{y}_{i})$ $\tilde{y}_{i}\in K_{i}\cap B_{\varepsilon}(\bar{y}_{i})$, $\widetilde{K}_{i}\in\Xi_{\varepsilon\rho}(\tilde{y}_{i})$ $(i=1,\ldots,n)$ and $x_{n+1}\in\Omega\cap B_{\varepsilon}(\bar{x})$, such that, for each $\hat{x}\in(\Omega-x_{n+1})\cap(\rho\mathbb{B}_{X})$, there is an $i\in\{1,\ldots,n\}$ such that
$F_{i}(\hat{x}+x_{i})\cap\big{(}y_{i}+(\widetilde{K}_{i}-\tilde{y}_{i})\cap(\rho\mathbb{B}_{Y})\big{)}=\emptyset.$
When $n=1$, Definition 4.1 reduces to Definition 3.2 (ii).
Application of Theorem 1.2 in the current setting leads to necessary conditions for approximate stationarity, and hence, stationarity and extremality extending Theorem 3.5 and its corollaries. Condition $({QC})_{C}$ can be extended as follows:
    1. $(\widetilde{QC})_{C}$
      there is an $\varepsilon>0$ such that $\big{\|}\sum_{i=1}^{n+1}x_{i}^{*}\big{\|}\geq\varepsilon$ for all $(x_{i},y_{i})\in{\rm gph}\,F_{i}\cap B_{\varepsilon}(\bar{x},\bar{y}_{i})$, $x_{i}^{*}\in D^{*C}F_{i}(x_{i},y_{i})(\varepsilon\mathbb{B}_{Y_{i}^{*}})$ $(i=1,\ldots,n)$, $x_{n+1}\in\Omega\cap B_{\varepsilon}(\bar{x})$ and $x_{n+1}^{*}\in N^{C}_{\Omega}(x_{n+1})$ such that $\sum_{i=1}^{n+1}\|x_{i}^{*}\|=1$,
while the corresponding extension $(\widetilde{QC})_{F}$ of condition $({QC})_{F}$ is obtained by replacing $N^{C}$ and $D^{*C}$ in $(\widetilde{QC})_{C}$ by $N^{F}$ and $D^{*F}$, respectively. An extension of Corollary 3.14 takes the following form.
Theorem 4.2.
Let $X$, $Y_{1},\ldots,Y_{n}$ be Banach spaces, $\Omega$ and, for each $i=1,\ldots,n$, the graph ${\rm gph}\,F_{i}$ and all members of $\Xi_{i}$ be closed. Suppose that the triple $\{F,\Omega,K\}$ is approximately stationary at $(\bar{x},\bar{y})$ with respect to $\Xi$. If condition $(\widetilde{QC})_{C}$ is satisfied, then there is an $M>0$ such that, for any $\varepsilon>0$, there exist $(x_{i},y_{i})\in{\rm gph}\,F_{i}\cap B_{\varepsilon}(\bar{x},\bar{y}_{i})$, $\widetilde{K}_{i}\in\Xi_{i}$, $v_{i}\in\widetilde{K}_{i}\cap B_{\varepsilon}(\bar{y}_{i})$, $y^{*}_{i}\in N^{C}_{\widetilde{K}_{i}}(v_{i})+\varepsilon\mathbb{B}_{Y_{i}^{*}}$ $(i=1,\ldots,n)$, and $x_{n+1}\in\Omega\cap B_{\varepsilon}(\bar{x})$ such that $\sum_{i=1}^{n}\|y_{i}^{*}\|=1$ and
$0\in\sum_{i=1}^{n}D^{*C}F_{i}(x_{i},y_{i})(y_{i}^{*})+N^{C}_{\Omega}(x_{n+1})\cap(M\mathbb{B}_{X^{*}})+\varepsilon\mathbb{B}_{X^{*}}.$
If $X$ is Asplund and condition $(\widetilde{QC})_{F}$ is satisfied, then the above assertion holds true with $N^{F}$ and $D^{*F}$ in place of $N^{C}$ and $D^{*C}$, respectively.
Remark 6.
  1. i.
    Proposition 3.8 (with $F:=(F_{1},\ldots,F_{n})$ in part (i)) gives two typical sufficient conditions for the fulfillment of conditions $(\widetilde{QC})_{C}$ and $(\widetilde{QC})_{F}$.
  2. ii.
    Theorem 4.2 covers [9, Theorems 3.1 and 3.2]. In view of the previous item, it also covers [9, Corollary 3.2].
  3. iii.
    Theorem 4.2 is a consequence of the dual necessary conditions for approximate stationarity of a collection of sets in Theorem 1.2. The latter theorem can be extended to cover a more general quantitative notion of approximate $\alpha$-stationarity, leading to corresponding extensions of Theorem 4.2 and its corollaries covering, in particular, dual conditions for $\varepsilon$-Pareto optimality in [21, Theorems 4.3 and 4.5].
Employing the multiple-mapping model studied in this section, one can consider a more general than ($P$) optimization problem with set-valued constraints:
$\text{minimize }\;F_{0}(x)\quad\text{subject to }\;F_{i}(x)\cap K_{i}\neq\emptyset\;(i=1,\ldots,n),\;\;x\in\Omega,$ ($\mathcal{P}$)
where $F_{i}:X\rightrightarrows Y_{i}$ $(i=0,\ldots,n)$ are mappings between normed spaces, $\Omega\subset X$, $K_{i}\subset Y_{i}$ $(i=1,\ldots,n)$, and $Y_{0}$ is equipped with a level-set mapping $L$. The “functional” constraints in ($\mathcal{P}$) can model a system of equalities and inequalities as well as more general operator-type constraints.
Using the set of admissible solutions $\widehat{\Omega}:=\{x\in\Omega\mid F_{i}(x)\cap K_{i}\neq\emptyset,\;i=1,\ldots,n\}$, we say that a point $(\bar{x},\bar{y}_{0})\in X\times Y_{0}$ is extremal in problem ($\mathcal{P}$) if it is extremal for $F_{0}$ on $\widehat{\Omega}$. This means, in particular, that $\bar{x}\in\Omega$, $\bar{y}_{0}\in F_{0}(\bar{x})$, and there exist $\bar{y}_{i}\in F_{i}(\bar{x})\cap K_{i}$ $(i=1,\ldots,n)$.
We are going to employ the model studied in the first part of this section with $n+1$ objects in place of $n$. There are $n+1$ mappings $F_{0},\ldots,F_{n}$ and $n$ sets $K_{1},\ldots,K_{n}$ in ($\mathcal{P}$). The missing set $K_{0}$ is determined by the level-set mapping $L$: we set $K_{0}:=L(\bar{y}_{0})\cup\{\bar{y}_{0}\}$. Given a $\delta>0$, we define the family of perturbations of $K_{0}$ by $\Xi_{0}^{\delta}:=\{{\rm cl}\,L(y)\mid y\in K_{0}\cap B_{\delta}(\bar{y}_{0})\}$. Now, set $Y:=Y_{0}\times\ldots\times Y_{n}$, $F:=(F_{0},\ldots,F_{n}):X\to Y$, $K:=K_{0}\times\ldots\times K_{n}$, $\bar{y}:=(\bar{y}_{0},\ldots,\bar{y}_{n})$ and $\Xi^{\delta}:=\Xi_{0}^{\delta}\times K_{1}\times\ldots\times K_{n}$. Observe from the last definition that only the $K_{0}$ component of $K$ is actually perturbed. Using the same arguments, one can prove the next extension of Proposition 3.16.
Proposition 4.3.
Let $\bar{x}\in\Omega$, $\bar{y}_{0}\in F_{0}(\bar{x})$, $\bar{y}_{i}\in F_{i}(\bar{x})\cap K_{i}$ $(i=1,\ldots,n)$, and $F$, $K$, $\bar{y}$ and $\Xi^{\delta}$ ${(\delta>0)}$ be defined as above. Suppose $L$ satisfies conditions (O1) and (O5). If $(\bar{x},\bar{y}_{0})$ is extremal in problem ($\mathcal{P}$), then $\{F,\Omega,K\}$ is extremal at $(\bar{x},\bar{y})$ with respect to $\Xi^{\delta}$ for all $\delta>0$.
Condition $(\widetilde{QC})_{C}$ in the current setting is reformulated as follows:
    1. $(\widetilde{QC})_{C}^{\prime}$
      there is an $\varepsilon>0$ such that $\big{\|}\sum_{i=0}^{n+1}x_{i}^{*}\big{\|}\geq\varepsilon$ for all $(x_{i},y_{i})\in{\rm gph}\,F_{i}\cap B_{\varepsilon}(\bar{x},\bar{y}_{i})$, $x_{i}^{*}\in D^{*C}F_{i}(x_{i},y_{i})(\varepsilon\mathbb{B}_{Y_{i}^{*}})$ $(i=0,\ldots,n)$, $x_{n+1}\in\Omega\cap B_{\varepsilon}(\bar{x})$ and $x_{n+1}^{*}\in N^{C}_{\Omega}(x_{n+1})$ such that $\sum_{i=0}^{n+1}\|x_{i}^{*}\|=1$,
while the corresponding reformulation $(\widetilde{QC})_{F}^{\prime}$ of condition $(\widetilde{QC})_{F}$ is obtained by replacing $N^{C}$ and $D^{*C}$ in $(\widetilde{QC})_{C}^{\prime}$ by $N^{F}$ and $D^{*F}$, respectively. In view of Proposition 4.3, Theorem 4.2 yields the following statement.
Corollary 4.4.
Let $X$, $Y_{0},\ldots,Y_{n}$ be Banach spaces, the sets $\Omega$, ${\rm gph}\,F_{i}$ $(i=0,\ldots,n)$ and $K_{i}$ $(i=1,\ldots,n)$ be closed, and $\delta>0$. Suppose $L$ satisfies conditions (O1) and (O5). If $(\bar{x},\bar{y}_{0})$ is extremal in problem ($\mathcal{P}$) and condition $(\widetilde{QC})_{C}^{\prime}$ is satisfied, then there is an $M>0$ such that, for any $\varepsilon>0$, there exist $(x_{i},y_{i})\in{\rm gph}\,F_{i}\cap B_{\varepsilon}(\bar{x},\bar{y}_{i})$ $(i=0,\ldots,n)$, $x_{n+1}\in\Omega\cap B_{\varepsilon}(\bar{x})$, $y\in K_{0}\cap B_{\delta}(\bar{y}_{0})$, $v_{0}\in{\rm cl}\,L(y)\cap B_{\varepsilon}(\bar{y}_{0})$, $y_{0}^{*}\in N^{C}_{{\rm cl}\,L(y)}(v_{0})+\varepsilon\mathbb{B}_{Y_{0}^{*}}$, $v_{i}\in K_{i}\cap B_{\varepsilon}(\bar{y}_{i})$ and $y_{i}^{*}\in N^{C}_{K_{i}}(v_{i})+\varepsilon\mathbb{B}_{Y_{i}^{*}}$ $(i=1,\ldots,n)$ such that $\sum_{i=0}^{n}\|y_{i}^{*}\|=1$ and
$0\in\sum_{i=0}^{n}D^{*C}F(x_{i},y_{i})(y_{i}^{*})+N^{C}_{\Omega}(x_{n+1})\cap(M\mathbb{B}_{X^{*}})+\varepsilon\mathbb{B}_{X^{*}}.$
If $X$ is Asplund and condition $(\widetilde{QC})_{F}^{\prime}$ is satisfied, then the above assertion holds true with $N^{F}$ and $D^{*F}$ in place of $N^{C}$ and $D^{*C}$, respectively.
Disclosure statement
No potential conflict of interest was reported by the authors.
Funding
Nguyen Duy Cuong is supported by Vietnam National Program for the Development of Mathematics 2021-2030 under grant number B2023-CTT-09.
Acknowledgments
A part of the work was done during Alexander Kruger’s stay at the Vietnam Institute for Advanced Study in Mathematics in Hanoi. He is grateful to the Institute for its hospitality and supportive environment.